Determine whether f is a function from Z to R if (a) f(n) = ±n (b) f(n) = √ n2 + 1 (c) f(n) = 1 n2−4


1
Expert's answer
2020-09-02T18:20:35-0400

(a). We have to check well defineness of f to be a function in domain and range set..

Now, for "n_1.n_2\in \mathbb{Z}" such that "n_1=n_2" , thus

Case-I: if f(n)=n then "f(n_1)=n_1=n_2=f(n_2)"

Case-II: if f(n)=-n, then "f(n_1)=-n_1=-n_2=f(n_2)"

Thus, both cases are well defined. Neveretheless, the function f should assign each n to a single value instead of two different values. Thus "f(n)=\pm n" is not a function.


(b) As, "f(n)=\sqrt{n^2+1}" , Now consider "n_1=n_2\in \mathbb{Z}"

Thus, "n_1^2=n_2^2\iff n_1^2+1=n_2^2+1\iff\sqrt{ n_1^2+1}=\sqrt{n_2^2+1}\iff f(n_1)=f(n_2)" "f(n_1)=\sqrt{n_1^2+1}" is well defined.

Hence, f is a function.

(c).Given "f(n)=n^2-4" , thus for "n_1=n_2\in \mathbb{Z}\implies f(n_1)=n_1^2-4=n_2^2-4=f(n_2)"

Thus, f is well defined and hence a function.